The height of a baseball hit into the air is given by the quadratic equation h(t)=-5(t-2.6)^2+35 , where time,t, is in seconds and height, h(t), is in metres.
a) What was the height of the ball when it was hit?
b)What is the maximum height of the ball?
c)Is the ball still in the air after 4s? Explain
d) When is the ball at a height of 24m?
Thanks sooo much :)

Answer by stanbon(75887)

The height of a baseball hit into the air is given by the quadratic equation h(t)=-5(t-2.6)^2+35 , where time,t, is in seconds and height, h(t), is in metres.
h(t) = -5(t^2 - 5.2t + 6.76) + 35
h(t) = -5t^2 + 26t - 33.8 + 35
h(t) = -5t^2 + 26t + 1.2
---------------------------
a) What was the height of the ball when it was hit?:: h(0) = 1.25 ft
-------
b)What is the maximum height of the ball?::
Max occurs when t = -b/(2a) = -26/(2*-5) = 2.6 seconds
Max height = h(2.6) = 35 ft
--------------------------------------
c)Is the ball still in the air after 4s? Explain
h(4) = 25.2 ft
d) When is the ball at a height of 24m?
Solve:: -5t^2 + 26t + 1.2 = 24
t = 1.12 sec and t = 4.08 sec
----------------
Cheers,
Stan H.
--------------